Verify the identity
1+sinx/cosx+cosx/1+sinx=2secx
1
Expert's answer
2011-03-09T07:41:55-0500
(1+sinx)/cosx+cosx/(1+sinx)=2secx

Let's consider the left part of the expression:
((1+sinx)2 + cos2x)/(cosx(1 +sinx)) = (1+2sinx +sin2x +cos2x)/(cosx*(1+sinx)) =

= (2 + 2 sinx)/(cosx*(1+sinx)) = 2(1+sinx)/(cosx (1+sinx)) = 2/cosx = 2 secx
2secx = 2secx

The identity is proved.